OpenStudy (anonymous):

What is the solution to the following system? -3x - y = -3 9x + 3y = 9 (3, 3) (3, -3) no solution infinite solutions on the line

OpenStudy (anonymous):

infinite

OpenStudy (anonymous):

yes the last option

OpenStudy (anonymous):

can u guys help me with this one

OpenStudy (anonymous):

to clear fractions from the equation 5/6n - 1/3n = -15 you could multiply each term on both sides by the number 6. True False

OpenStudy (anonymous):

yes true

OpenStudy (anonymous):

If a line has a slope of 2/5 then a perpendicular line would have a slope of-5/2 true false

OpenStudy (anonymous):

true

OpenStudy (anonymous):

If you were asked to define 3 consecutive odd integers, you would name them x, x + 2, x + 4. True False

OpenStudy (anonymous):

true

OpenStudy (anonymous):

infinite

OpenStudy (anonymous):

The polynomial 15xy+3x3y4 is what degree? 6 5 7 9

OpenStudy (anonymous):

5

OpenStudy (anonymous):

The graph of the solution to the inequality -5x – 12 > 3 has an open circle at -3 and is shaded to the left. True False

OpenStudy (anonymous):

??

OpenStudy (anonymous):

true true true

OpenStudy (anonymous):

What is the slope of the line for the equation y = 3? 3 0 undefined none of the above

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Solve the absolute value equation |2n + 15| = 1. One solution is n = -7. What is the other solution? n = -14 n = 8 n = 7 none of the above

OpenStudy (anonymous):

??

OpenStudy (anonymous):

none

OpenStudy (sirm3d):

the degree of the polynomial 15xy +3x3y4 is 7, not 5.

OpenStudy (anonymous):

thats the power i thought that was coefficient so its 7 you are right

OpenStudy (sirm3d):

the graph of the ... is an open circle ... TRUE

OpenStudy (anonymous):

What is the slope of the line 2x – 5y = -10

OpenStudy (sirm3d):

2/5

OpenStudy (anonymous):

Identify the x and y intercepts for the equation 4x – 3y = -12. (4, 0) and (0, -3) (3, 0) and (0, -4) (4, 0) and (0, 3) (-3, 0) and (0, 4) (3, 0) and (0, 4)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(-3, 0) and (0, 4)

OpenStudy (sirm3d):

the fourth one, (-3,0) and (0,4)

OpenStudy (anonymous):

how good are u at this math stuff ?

OpenStudy (anonymous):

What type of polynomial is 3x2+6x+1? monomial binomial trinomial not a polynomial

OpenStudy (sirm3d):

do we have to answer the other question? LOL

OpenStudy (sirm3d):

trinomial

OpenStudy (anonymous):

lol why

OpenStudy (sirm3d):

next question, pls. hihi